The starting salaries of individuals with an MBA degree are normally distributed with a mean of $65,000 and a standard deviation of $15,000. What is the lowest salary for those individuals with an MBA degree whose starting salary is in the top 40 percent? Explain how you got the right answer

Answers: a.

$66,400

b.

$63,750

c.

$68,750

d.

$72,600
Answer #1

Solution:-

Given that,

mean = = 65000

standard deviation = =15000

Using standard normal table,

P(Z > z) = 40%

= 1 - P(Z < z) = 0.40  

= P(Z < z) = 1 - 0.40

= P(Z < z ) = 0.60

= P(Z < 0.25) = 0.60

z =0.25 ( using z table )

Using z-score formula,

x = z * +

x =0.25 * 15000+65000

x = 68750